OpenStudy (anonymous):

Use E = F/q to solve the following problem: Ernie pulls his wool sweater over his head, which charges his body as the sweater rubs against his cotton shirt. What is the electric field at a location where a 1.60 x 10^-19 C piece of lint experiences a force of 3.2 x 10^-9 N as it floats near Ernie?

OpenStudy (unklerhaukus):

\[\vec E=\frac{ \vec F}C\]\[\vec E=\frac{3.2\times10^{-9}\text N }{1.60\times 10^{-9} \text C }\]

OpenStudy (anonymous):

\[E=F/Q\] E=(3.2*10^-9)/(1.60*10^-19) =2*10^10 volt/metre or newton/coulomb